# Many-body Fu-Kane-Mele index

### v1

@@ 5.2.1 Flux insertion for spin Hall states in the free fermion formalism
see [4].
Let ${\mathscr{N}}_{\pm}$ be the orthogonal projections onto $\mathrm{Ker}({\mathscr{P}}_{+}^{\uparrow}-{\mathscr{P}}_{-}^{\uparrow}\mp\operatorname{\mathbbm{1}})$ and let $n_{\pm}=\dim{\mathscr{N}}_{\pm}$ be the dimensions of these kernels. With this, (5.24) reads $n_{+}-n_{-}=\mathrm{Ch}({\mathscr{P}}^{\uparrow})$. We define
${\mathscr{Q}}_{\pm}^{\uparrow}:={\mathscr{P}}_{\pm}^{\uparrow}-{\mathscr{N}}_{\pm}$ (5.26)
and note that
Lemma 5.4.
The operators ${\mathscr{Q}}_{\pm}^{\uparrow}$ are projections and ${\mathscr{Q}}_{+}^{\uparrow}-{\mathscr{Q}}_{-}^{\uparrow}\in{\mathcal{J}}_{1}$. Moreover, $\mathrm{ind}({\mathscr{Q}}_{+}^{\uparrow},{\mathscr{Q}}_{-}^{\uparrow})=0$.
Proof.
We first show that the projection ${\mathscr{N}}_{+}$ is a subprojection of ${\mathscr{P}}_{+}^{\uparrow}$, and similarly for ${\mathscr{N}}_{-}$ and ${\mathscr{P}}_{-}^{\uparrow}$. i.e. we want to show that ${\mathscr{N}}_{\pm}{\mathscr{P}}_{\pm}^{\uparrow}{\mathscr{N}}_{\pm}={\mathscr{N}}_{\pm}$. First, $0\leq{\mathscr{N}}_{\pm}{\mathscr{P}}_{\pm}^{\uparrow}{\mathscr{N}}_{\pm}\leq{\mathscr{N}}_{\pm}$ follows from $0\leq{\mathscr{P}}_{\pm}^{\uparrow}\leq\operatorname{\mathbbm{1}}$ and the fact that ${\mathscr{N}}_{\pm}$ is self-adjoint. Second, the definition of ${\mathscr{N}}_{\pm}$ is equivalent to ${\mathscr{N}}_{\pm}({\mathscr{P}}_{+}^{\uparrow}-{\mathscr{P}}_{-}^{\uparrow}){\mathscr{N}}_{\pm}=\pm{\mathscr{N}}_{\pm}$, which implies that
${\mathscr{N}}_{\pm}{\mathscr{P}}_{\pm}^{\uparrow}{\mathscr{N}}_{\pm}={\mathscr{N}}_{\pm}(\operatorname{\mathbbm{1}}+{\mathscr{P}}_{\mp}^{\uparrow}){\mathscr{N}}_{\pm}\geq{\mathscr{N}}_{\pm}.$ (5.27)
We conclude that ${\mathscr{N}}_{\pm}{\mathscr{P}}_{\pm}={\mathscr{P}}_{\pm}{\mathscr{N}}_{\pm}={\mathscr{N}}_{\pm}$. It follow immediately that ${\mathscr{Q}}_{\pm}^{\uparrow}={\mathscr{P}}_{\pm}^{\uparrow}-{\mathscr{N}}_{\pm}$ is a projection. Since ${\mathscr{N}}_{\pm}$ are finite range projections the claim that ${\mathscr{Q}}_{+}^{\uparrow}-{\mathscr{Q}}_{-}^{\uparrow}$ is trace class follows immediately from the fact that ${\mathscr{P}}_{+}^{\uparrow}-{\mathscr{P}}_{-}^{\uparrow}$ is trace class (Proposition B.6). It follows that $\mathrm{ind}({\mathscr{Q}}_{+}^{\uparrow},{\mathscr{Q}}_{-}^{\uparrow})$ is well-defined and we compute it as follows:
$\mathrm{ind}({\mathscr{Q}}_{\pm}^{\uparrow},{\mathscr{P}}_{\pm}^{\uparrow})=\mathrm{dim}\mathrm{Ker}(-{\mathscr{N}}_{\pm}-\operatorname{\mathbbm{1}})-\mathrm{dim}\mathrm{Ker}(-{\mathscr{N}}_{\pm}+\operatorname{\mathbbm{1}})=0-\mathrm{dim}{\mathscr{N}}_{\pm}.$ (5.28)
Hence, by the additivity of the index of a pair of projections  [4, Theorem 3.4],
$\mathrm{ind}({\mathscr{Q}}_{+}^{\uparrow},{\mathscr{Q}}_{-}^{\uparrow})$ $=\mathrm{ind}({\mathscr{Q}}_{+}^{\uparrow},{\mathscr{P}}_{+}^{\uparrow})+\mathrm{ind}({\mathscr{P}}_{+}^{\uparrow},{\mathscr{P}}_{-}^{\uparrow})-\mathrm{ind}({\mathscr{Q}}_{-}^{\uparrow},{\mathscr{P}}_{-}^{\uparrow})$ (5.29)
$=-n_{+}+\mathrm{Ch}({\mathscr{P}}^{\uparrow})+n_{-}=0$ (5.30)
where we used $\mathrm{ind}({\mathscr{P}}_{+}^{\uparrow},{\mathscr{P}}_{-}^{\uparrow})=\mathrm{Ch}({\mathscr{P}}^{\uparrow})=n_{+}-n_{-}$. ∎
Since ${\mathscr{Q}}_{+}^{\uparrow}-{\mathscr{Q}}_{-}^{\uparrow}\in{\mathcal{J}}_{1}$ we can apply to the following Proposition:
Proposition 5.5.
Let ${\mathscr{P}},{\mathscr{Q}}$ be projections acting on a Hilbert space ${\mathcal{H}}$ with ${\mathscr{P}}-{\mathscr{Q}}\in{\mathcal{J}}_{1}$ and $\mathrm{ind}({\mathscr{P}},{\mathscr{Q}})=0$. Then there exists a unitary ${\mathscr{V}}$ such that ${\mathscr{P}}={\mathscr{V}}{\mathscr{Q}}{\mathscr{V}}^{*}$ and ${\mathscr{V}}-\operatorname{\mathbbm{1}}\in{\mathcal{J}}_{1}$.
The proof can be found in Section VII B of [15] (without the trace class statement) or in Section 4.2.1 of [11] (including the trace class statement).
From Lemma 5.4 and Proposition 5.5 we get a unitary ${\mathscr{V}}$ such that
${\mathscr{Q}}_{+}^{\uparrow}={\mathscr{V}}{\mathscr{Q}}_{-}^{\uparrow}{\mathscr{V}}^{*},\qquad{\mathscr{V}}-\operatorname{\mathbbm{1}}\in{\mathcal{J}}_{1}.$ (5.31)
Using this and (5.23) we get for ${\mathscr{P}}_{+}={\mathscr{P}}_{+}^{\uparrow}\oplus{\mathscr{P}}_{+}^{\downarrow}\oplus 0$ that
${\mathscr{P}}_{+}$ $=({\mathscr{Q}}_{+}^{\uparrow}+{\mathscr{N}}_{+})\oplus({\mathscr{K}}{\mathscr{P}}_{-}^{\uparrow}{\mathscr{K}})\oplus 0$ (5.32)
$=({\mathscr{V}}{\mathscr{Q}}_{-}^{\uparrow}{\mathscr{V}}^{*}+{\mathscr{N}}_{+})\oplus{\mathscr{K}}({\mathscr{Q}}_{-}^{\uparrow}+{\mathscr{N}}_{-}){\mathscr{K}}\oplus 0$ (5.33)
$=({\mathscr{V}}\oplus\operatorname{\mathbbm{1}}\oplus\operatorname{\mathbbm{1}})({\mathscr{Q}}_{-}^{\uparrow}\oplus{\mathscr{K}}{\mathscr{Q}}_{-}^{\uparrow}{\mathscr{K}}\oplus 0)({\mathscr{V}}\oplus\operatorname{\mathbbm{1}}\oplus\operatorname{\mathbbm{1}})^{*}+({\mathscr{N}}_{+}\oplus({\mathscr{K}}{\mathscr{N}}_{-}{\mathscr{K}})\oplus 0).$ (5.34)
The projections ${\mathscr{P}}_{+}$ and $({\mathscr{V}}\oplus\operatorname{\mathbbm{1}}\oplus\operatorname{\mathbbm{1}})({\mathscr{Q}}_{-}^{\uparrow}\oplus{\mathscr{K}}{\mathscr{Q}}_{-}^{\uparrow}{\mathscr{K}}\oplus 0)({\mathscr{V}}\oplus\operatorname{\mathbbm{1}}\oplus\operatorname{\mathbbm{1}})^{*}$ differ by a finite rank operator and hence the quasi-free states associated with ${\mathscr{P}}_{+}$ and ${\mathscr{Q}}_{-}^{\uparrow}\oplus{\mathscr{K}}{\mathscr{Q}}_{-}^{\uparrow}{\mathscr{K}}\oplus 0$ are unitarily equivalent. Moreover,
$[({\mathscr{V}}\oplus\operatorname{\mathbbm{1}}\oplus\operatorname{\mathbbm{1}}),({\mathscr{Q}}_{-}^{\uparrow}\oplus{\mathscr{K}}{\mathscr{Q}}_{-}^{\uparrow}{\mathscr{K}}\oplus 0)]=[{\mathscr{V}},{\mathscr{Q}}_{-}^{\uparrow}]\in{\mathcal{J}}_{1}$ (5.35)
because ${\mathscr{V}}-\operatorname{\mathbbm{1}}\in{\mathcal{J}}_{1}$. The Shale-Stinespring criterion [50], see also [20, Theorem 6.16], implies that the Bogoliubov transformation $a(f)\mapsto a(({\mathscr{V}}\oplus\operatorname{\mathbbm{1}}\oplus\operatorname{\mathbbm{1}})f)$ is unitarily implementable in the GNS representation of the quasi-free state associated with the projection ${\mathscr{Q}}_{-}^{\uparrow}\oplus{\mathscr{K}}{\mathscr{Q}}_{-}^{\uparrow}{\mathscr{K}}\oplus 0$. We denote the corresponding unitary by $\Gamma({\mathscr{V}}\oplus\operatorname{\mathbbm{1}}\oplus\operatorname{\mathbbm{1}})$.
To the projection ${\mathscr{R}}={\mathscr{Q}}_{-}^{\uparrow}\oplus{\mathscr{K}}{\mathscr{Q}}_{-}^{\uparrow}{\mathscr{K}}\oplus 0$ corresponds a quasi-free state $\omega_{{\mathscr{R}}}$ whose GNS triple we denote by $(\Pi,{\mathcal{H}},|\Phi\rangle)$. One then checks using the definition (2.10) of ${\mathscr{T}}$ that ${\mathscr{T}}({\mathscr{Q}}_{-}^{\uparrow}\oplus{\mathscr{K}}{\mathscr{Q}}_{-}^{\uparrow}{\mathscr{K}}\oplus 0){\mathscr{T}}^{*}={\mathscr{Q}}_{-}^{\uparrow}\oplus{\mathscr{K}}{\mathscr{Q}}_{-}^{\uparrow}{\mathscr{K}}\oplus 0$, so that $\omega_{\mathscr{R}}$ is time reversal invariant. It follows that the time reversal $\tau$ induced by the single particle time reversal ${\mathscr{T}}$ is implemented in the GNS representation by an anti-unitary operator $T$ which we may choose such that
$T|\Phi\rangle=|\Phi\rangle$ (5.36)
and therefore
$T^{2}|\Phi\rangle=|\Phi\rangle.$ (5.37)
We see that $|\Phi\rangle$ is a Kramers singlet. By the remarks above, the quasi-free state corresponding to ${\mathscr{P}}_{+}$ is represented by a vector $|\Psi_{+}\rangle\in{\mathcal{H}}$ given by
$|\Psi_{+}\rangle:=A^{*}\Gamma({\mathscr{V}}\oplus\operatorname{\mathbbm{1}}\oplus\operatorname{\mathbbm{1}})|\Phi\rangle$ (5.38)
where
$A^{*}=\Pi\Big{(}\prod_{i=1}^{n_{+}}a_{\uparrow}^{*}(v_{i})\prod_{j=1}^{n_{-}}a_{\downarrow}^{*}({\mathscr{K}}w_{i})\Big{)}.$ (5.39)
Here $\{v_{i}\}_{i=1}^{n_{+}}$ and $\{w_{i}\}_{i=1}^{n_{-}}$ are orthonormal bases of $\mathrm{Ran}{\mathscr{N}}_{+}$ and $\mathrm{Ran}{\mathscr{N}}_{-}$ respectively. Being a product of creation operators, $A^{*}$ is homogeneous with
$T^{2}A^{*}T^{2}=(-1)^{n_{+}+n_{-}}A^{*}$ (5.40)
since $T^{2}$ implements fermion parity. Since $(-1)^{n_{+}+n_{-}}=(-1)^{n_{+}-n_{-}}$ and $n_{+}-n_{-}$ equals the Chern number of ${\mathscr{P}}^{\uparrow}$ we conclude that
$T^{2}|\Psi_{+}\rangle=(-1)^{\mathrm{Ch}({\mathscr{P}}^{\uparrow})}A^{*}T^{2}\Gamma({\mathscr{V}}\oplus\operatorname{\mathbbm{1}}\oplus\operatorname{\mathbbm{1}})|\Phi\rangle=(-1)^{\mathrm{Ch}({\mathscr{P}}^{\uparrow})}|\Psi_{+}\rangle$ (5.41)
where we used that $\Gamma({\mathscr{V}}\oplus\operatorname{\mathbbm{1}}\oplus\operatorname{\mathbbm{1}})|\Phi\rangle$ is even under $T^{2}$ because $\Gamma({\mathscr{V}}\oplus\operatorname{\mathbbm{1}}\oplus\operatorname{\mathbbm{1}})$ is an even unitary and $|\Phi\rangle$ is a Kramers singlet, see [44] and (5.37). This shows that the quasi-free state $\omega_{{\mathscr{P}}_{+}}$ corresponding to ${\mathscr{P}}_{+}$ is a Kramers singlet if $\mathrm{Ch}({\mathscr{P}}^{\uparrow})$ is even, while it belongs to a Kramers pair if $\mathrm{Ch}({\mathscr{P}}^{\uparrow})$ is odd.
5.2.2 Non-triviality of the index for spin Hall states
@@
The state $\psi^{+}:=\omega_{{\mathscr{P}}_{+}}$ was obtained from $\omega=\omega_{\mathrm{SH}}\,\hat{\otimes}\,\omega_{\mathrm{empty}}$ by inserting a $\pi$ flux defect using a Kato generator (5.18) in the single particle Hilbert space. Since we defined our index using defect states obtained through a quasi-adiabatic evolution rather than Kato’s parallel transport, it remains to show that both flux insertion protocols lead to the same index. The argument follows the strategy used in the proof of Proposition 3.8 to show that $\omega^{+}$ and $\psi^{+}$ have the same parity.
@@
In order to compare the defect state $\psi^{+}$ obtained using the Kato generator with a defect state $\omega^{+}$ obtained by quasi-adiabatic evolution we first give a many-body description of the Kato generator (5.18). It is shown in the Appendix, Lemma B.1, that ${\mathscr{K}}_{\phi}^{h}$ is exponentially local, and exponentially localised near the horizontal axis. i.e. writing
${\mathscr{K}}^{h}_{\phi}(x,i;y,j):=\langle\delta_{x}\otimes e_{i},{\mathscr{K}}^{h}_{\phi}\delta_{y}\otimes e_{j}\rangle$ (5.42)
we have
@@
for some positive constants $c$ and $\xi$. It follows immediately that the Kato generator ${\mathscr{K}}_{\phi}$, see (5.18), is exponentially localised near the left horizontal axis $l=(0,-(1,0))$.
In particular, ${\mathscr{K}}_{\phi}$ satisfies the assumptions of Lemma 5.3 so that the projections ${\mathscr{P}}_{\phi}$ defined in (5.20) correspond to quasi-free pure states $\psi^{\phi}$ such that $\psi^{+}=\psi^{\pi}$ and there is a TDI $K^{l}$ such that
@@
for all $\phi\in[-\pi,\pi]$. The TDI may be taken to be (see the proof of Lemma 5.3)
@@
There is an $f\in{\mathcal{F}}$ for each $\phi\in[0,\pi]$ an $f$-localized even unitary $V_{\phi}\in{\mathcal{A}}^{\mathrm{al}}$ such that $\omega^{\phi}=\psi^{\phi}\circ\mathrm{Ad}[V_{\phi}]$.
@@
Since $\omega$ is SRE there is a homogeneous product state $\omega_{0}$ and an LGA $\alpha$ such that $\omega=\omega_{0}\circ\alpha$. For all $\phi\in{\mathbb{R}}$ we now put
$\tilde{\psi}^{\phi}$ $=\psi^{\phi}\circ(\alpha\circ\alpha_{\phi}^{K})^{-1}$ (5.48)
$\tilde{\omega}^{\phi}$ $=\omega^{\phi}\circ(\alpha\circ\alpha_{\phi}^{K})^{-1},$ (5.49)
@@
$\tilde{\beta}^{\phi}=\alpha\circ\gamma_{\phi}^{l,H}\circ(\alpha_{\phi}^{K})^{-1}\circ\alpha^{-1}.$ (5.51)
@@
Arguing as in the first part of the proof of Lemma 3.7, we can find $g\in{\mathcal{F}}$ uniform in $\phi\in[0,\pi]$ such that
$\|\tilde{\beta}^{\phi}(A)-\beta^{\phi}(A)\|\leq g(r)\|A\|$ (5.53)
@@
$\left|\tilde{\psi}^{\phi}(A)-\tilde{\omega}^{\phi}(A)\right|=\left|\omega_{0}\big{(}\beta^{\phi}(A)-\tilde{\beta}^{\phi}(A)\big{)}\right|\leq g(r)\left|A\right|$ (5.54)
for all $A\in\Lambda_{\pi,l}\cap B(r)^{c}$. Similarly we have $\|\tilde{\beta}(A)-A\|\leq g(r)\|A\|$ for $A\in\Lambda_{\pi,-l}\cap B(r)^{c}$ which implies $\left|\tilde{\psi}^{\phi}(A)-\tilde{\omega}^{\phi}(A)\right|\leq g(r)\|A\|$ for $A\in\Lambda_{\pi,l}\cap B(r)^{c}$.
Since $\tilde{\psi}^{\phi}=\omega_{0}$ is a homogeneous product state it now follows from Lemma A.1 that the states $\tilde{\omega}^{\phi}$ are all $h$-close to $\omega_{0}$ for $\phi\in[0,\pi]$ for an $h\in{\mathcal{F}}$ that depends only on $g$, and by Proposition A.2 there is an $\tilde{f}\in{\mathcal{F}}$ depending only on $h$, and $\tilde{f}$-local homogeneous unitaries $W_{\phi}\in{\mathcal{A}}^{\mathrm{al}}$ such that $\tilde{\omega}^{\phi}=\omega_{0}\circ\mathrm{Ad}[W_{\phi}]$. It follows that $\omega^{\phi}=\psi^{\phi}\circ\mathrm{Ad}[V_{\phi}]$ with $V_{\phi}=(\alpha\circ\alpha_{\phi}^{K})^{-1}(W_{\phi})$, which are uniformly $f$-local for some $f\in{\mathcal{F}}$.
@@
$\mathrm{Ind}_{2}(\omega_{\mathrm{SH}})=(-1)^{\mathrm{Ch}({\mathscr{P}}^{\uparrow})}$.

### v2

@@ 5.2 Does the state $\omega_{{\mathscr{Q}}^{\mathrm{qa}}_{\pi}}$ belong to a Kramers pair?
see [7]. Since the difference ${\mathscr{Q}}^{\mathrm{qa}}_{\pi}-{\mathscr{P}}_{\pi}^{(\mu)}$ is trace class by Proposition 5.1 the index of the pair $({\mathscr{Q}}^{\mathrm{qa}}_{\pi},{\mathscr{P}}_{\pi}^{(\mu)})$ is well-defined and given by the difference of the ranks of the orthogonal projectors ${\mathscr{N}}_{\pm}$ onto the finite dimensional kernels of ${\mathscr{Q}}^{\mathrm{qa}}_{\pi}-{\mathscr{P}}_{\pi}^{(\mu)}\mp\operatorname{\mathbbm{1}}$. We claim that ${\mathscr{N}}_{+}$ is a subprojector of ${\mathscr{Q}}^{\mathrm{qa}}_{\pi}$ while ${\mathscr{N}}_{-}$ is orthogonal to ${\mathscr{Q}}^{\mathrm{qa}}_{\pi}$. First, $0\leq{\mathscr{N}}_{\pm}{\mathscr{Q}}^{\mathrm{qa}}_{\pi}{\mathscr{N}}_{\pm}\leq{\mathscr{N}}_{\pm}$ follows from $0\leq{\mathscr{Q}}^{\mathrm{qa}}_{\pi}\leq\operatorname{\mathbbm{1}}$ and the fact that ${\mathscr{N}}_{\pm}$ is an orthogonal projector. Second, the definition of ${\mathscr{N}}_{\pm}$ is equivalent to ${\mathscr{N}}_{\pm}({\mathscr{Q}}^{\mathrm{qa}}_{\pi}-{\mathscr{P}}_{\pi}^{(\mu)}){\mathscr{N}}_{\pm}=\pm{\mathscr{N}}_{\pm}$, which implies that
${\mathscr{N}}_{\pm}{\mathscr{Q}}^{\mathrm{qa}}_{\pi}{\mathscr{N}}_{\pm}={\mathscr{N}}_{\pm}(\pm\operatorname{\mathbbm{1}}+{\mathscr{P}}_{\pi}^{(\mu)}){\mathscr{N}}_{\pm}.$ (5.17)
Hence ${\mathscr{N}}_{+}{\mathscr{Q}}^{\mathrm{qa}}_{\pi}{\mathscr{N}}_{+}\geq{\mathscr{N}}_{+}$. We conclude that ${\mathscr{N}}_{+}{\mathscr{Q}}^{\mathrm{qa}}_{\pi}={\mathscr{Q}}^{\mathrm{qa}}_{\pi}{\mathscr{N}}_{+}={\mathscr{N}}_{+}$ and in turn that ${\mathscr{P}}_{\pi}^{(\mu)}{\mathscr{N}}_{+}={\mathscr{N}}_{+}{\mathscr{P}}_{\pi}^{(\mu)}=0$. The second alternative of (5.17) yields ${\mathscr{N}}_{-}{\mathscr{Q}}^{\mathrm{qa}}_{\pi}={\mathscr{Q}}^{\mathrm{qa}}_{\pi}{\mathscr{N}}_{-}=0$. Since, moreover, ${\mathscr{N}}_{\pm}$ are orthogonal to each other, we conclude that $\widetilde{\mathscr{Q}}^{\mathrm{qa}}_{\pi}:={\mathscr{Q}}^{\mathrm{qa}}_{\pi}-{\mathscr{N}}_{+}+{\mathscr{N}}_{-}$ is an orthogonal projector. Hence,
$\mathrm{ind}(\widetilde{\mathscr{Q}}^{\mathrm{qa}}_{\pi},{\mathscr{P}}_{\pi}^{(\mu)})=\operatorname{Tr}\{\widetilde{\mathscr{Q}}^{\mathrm{qa}}_{\pi}-{\mathscr{P}}_{\pi}^{(\mu)}\}=\operatorname{Tr}\{{\mathscr{Q}}^{\mathrm{qa}}_{\pi}-{\mathscr{P}}_{\pi}^{(\mu)}\}-\operatorname{Tr}\{{\mathscr{N}}_{+}-{\mathscr{N}}_{-}\}=0,$ (5.18)
where we used the fact that the index of a pair of projections whose difference is trace class is given by the trace of their difference (cf. Proposition 2.2 of [6])). It now follows from the argument in Section 4.2.1 of [14] (see also Section VII B of [18]) that there exists a unitary ${\mathscr{V}}$ such that ${\mathscr{V}}-\operatorname{\mathbbm{1}}$ is trace class and $\widetilde{\mathscr{Q}}^{\mathrm{qa}}_{\pi}={\mathscr{V}}{\mathscr{P}}_{\pi}^{(\mu)}{\mathscr{V}}^{*}$. We therefore have
${\mathscr{Q}}^{\mathrm{qa}}_{\pi}={\mathscr{V}}{\mathscr{P}}_{\pi}^{(\mu)}{\mathscr{V}}^{*}-{\mathscr{N}}_{+}+{\mathscr{N}}_{-}.$ (5.19)
Since ${\mathscr{H}}_{-\pi}={\mathscr{H}}_{\pi}$ we have ${\mathscr{T}}{\mathscr{P}}^{(\mu)}_{\pi}{\mathscr{T}}^{*}={\mathscr{P}}_{-\pi}^{(\mu)}={\mathscr{P}}_{\pi}^{(\mu)}$ is time reversal invariant and therefore so is the quasi-free state $\omega_{{\mathscr{P}}_{\pi}^{(\mu)}}$. By Lemma 3.9 there is an antiunitary $T$ that implements time reversal in the GNS representation $(\Pi_{\pi},{\mathcal{H}}_{\pi},\Omega)$ of $\omega_{{\mathscr{P}}_{\pi}^{(\mu)}}$ which we can take such that $T\Omega=\Omega$. It follows from (5.19) that the quasi-free state $\omega_{{\mathscr{Q}}^{\mathrm{qa}}_{\pi}}$ is represented in ${\mathcal{H}}_{\pi}$ by the vector
$\Omega^{+}=AB\Gamma({\mathscr{V}})\Omega$ (5.20)
where $\Gamma({\mathscr{V}})$ is the unitary which implements the Bogoliubov transformation $a(f)\mapsto a({\mathscr{V}}f)$ (existence of $\Gamma({\mathscr{V}})$ is guaranteed by the Shale-Stinespring criterion [50], see also [21, Theorem 6.16]) and
$A=\Pi_{\pi}\left(\prod_{i=1}^{\dim{\mathscr{N}}_{-}}\,a^{*}(u_{i})\right),\quad B=\Pi_{\pi}\left(\prod_{i=1}^{\dim{\mathscr{N}}_{+}}\,a(v_{i})\right)$ (5.21)
where $\{u_{i}\}$ and $\{v_{i}\}$ are orthonormal bases of the range of ${\mathscr{N}}_{-}$ and ${\mathscr{N}}_{+}$ respectively. Since $\Gamma({\mathscr{V}})$ is even (it preserves particle number) we find
$T^{2}\Omega^{+}$ $=T^{*}AB\Gamma({\mathscr{V}})\Omega=(-1)^{\dim{\mathscr{N}}_{+}-\dim{\mathscr{N}}_{-}}\,\Omega^{+}$ (5.22)
$=(-1)^{\operatorname{Tr}\{{\mathscr{Q}}^{\mathrm{qa}}_{\pi}-{\mathscr{P}}_{\pi}^{(\mu)}\}}\,\Omega^{+}=(-1)^{\operatorname{\mathrm{ind}_{\mathrm{FKM}}}({\mathscr{P}})}\,\Omega^{+}$ (5.23)
where we used Proposition 5.1 in the last equality. This concludes the proof. ∎
5.3 The index extends the Fu-Kane-Mele index for $\operatorname{\mathrm{AII}}$ states
@@
The state $\psi^{+}:=\omega_{{\mathscr{Q}}^{\mathrm{qa}}_{\pi}}$ was obtained from $\omega:=\omega_{{\mathscr{P}}}$ by inserting a $\pi$ flux defect using a quasi-adiabatic generator (5.5) in the single particle Hilbert space. Since we defined our many-body index using defect states obtained through Hasting’s quasi-adiabatic evolution, it remains to show that both flux insertion protocols lead to the same index. For this, we show that states obtained using the second quantization of (5.5) and using Hastings’ evolution are unitarily equivalent for all values of the flux $\phi\in[0,\pi]$. The argument follows the same strategy used in the proof of Proposition 3.8.
@@
In order to compare the defect state $\psi^{+}$ obtained using the free fermion generator with a defect state $\omega^{+}$ obtained by Hasting’s quasi-adiabatic evolution we first give a many-body description of the free fermion generator (5.5). It is shown in the Appendix, Lemma C.1, that ${\mathscr{K}}_{\phi}^{h}$ is exponentially local, and exponentially localised near the horizontal axis, namely
@@
for some positive constants $c$ and $\xi$. It follows immediately that the free fermion generator ${\mathscr{K}}_{\phi}$, see (5.5), is exponentially localised near the left horizontal axis $l=(0,-(1,0))$.
In particular, ${\mathscr{K}}_{\phi}$ satisfies the assumptions of Lemma B.3 so that the projections ${\mathscr{Q}}^{\mathrm{qa}}_{\phi}$ defined in (5.7) correspond to quasi-free pure states $\psi^{\phi}$ such that $\psi^{+}=\psi^{\pi}$ and there is a TDI $K^{l}$ such that
@@
for all $\phi\in[-\pi,\pi]$. The TDI may be taken to be (see the proof of Lemma B.3)
@@
There is an $f\in{\mathcal{F}}$ and for each $\phi\in[0,\pi]$ an $f$-localized even unitary $V_{\phi}\in{\mathcal{A}}^{\mathrm{al}}$ such that $\omega^{\phi}=\psi^{\phi}\circ\mathrm{Ad}[V_{\phi}]$.
@@
Since $\omega$ is SRE there is a homogeneous product state $\omega_{0}$ and an LGA $\alpha$ such that $\omega=\omega_{0}\circ\alpha$. For all $\phi\in{\mathbb{R}}$ we define
$\tilde{\psi}^{\phi}=\psi^{\phi}\circ(\alpha\circ\alpha_{\phi}^{K^{l}})^{-1},\quad\text{and}\quad\tilde{\omega}^{\phi}=\omega^{\phi}\circ(\alpha\circ\alpha_{\phi}^{K^{l}})^{-1}.$ (5.29)
@@
$\tilde{\beta}^{\phi}=\alpha\circ\gamma_{\phi}^{l,H}\circ(\alpha_{\phi}^{K^{l}})^{-1}\circ\alpha^{-1}.$ (5.31)
@@
We now argue by locality again. Because $K^{l}$ is a restriction of $K^{h}$ to the left half plane, see (5.5) and (5.26,5.27), the LGAs $\alpha_{\phi}^{K^{l}}$ and $\alpha_{\phi}^{K^{h}}$ are almost equal on any observable localized in the left half-plane $A\in\Lambda_{\pi,l}\cap B(r)^{c}$, see Lemma 3.5. The same holds for $\gamma_{\phi}^{l,H}$ and $\alpha_{\phi}^{h,H}$ with the former being generated by the restriction of the generator of the latter to the left half-plane. We conclude that there is $g_{1}\in{\mathcal{F}}$ such that
$\|\tilde{\beta}^{\phi}(A)-\beta^{\phi}(A)\|\leq g_{1}(r)\|A\|$ (5.33)
@@
$\left|\tilde{\psi}^{\phi}(A)-\tilde{\omega}^{\phi}(A)\right|=\left|\omega_{0}\big{(}\beta^{\phi}(A)-\tilde{\beta}^{\phi}(A)\big{)}\right|\leq g_{1}(r)\|A\|$ (5.34)
for all $A\in\Lambda_{\pi,l}\cap B(r)^{c}$. What is more, both $\gamma_{\phi}^{l,H}$ and $\alpha_{\phi}^{K^{l}}$ act almost trivially on any observable localized in the right half-plane $A\in\Lambda_{\pi,l}\cap B(r)^{c}$, again by Lemma 3.5 and the fact that the generators are almost localised on the left half-plane. In turn, this implies that there is $g_{2}\in{\mathcal{F}}$ such that
$\|\tilde{\beta}(A)-A\|\leq g_{2}(r)\|A\|$ (5.35)
for all $\phi\in[0,\pi]$ and all $A\in\Lambda_{\pi,-l}\cap B(r)^{c}$. This implies that
$\left|\tilde{\psi}^{\phi}(A)-\tilde{\omega}^{\phi}(A)\right|=\left|\omega_{0}\big{(}A-\tilde{\beta}^{\phi}(A)\big{)}\right|\leq g_{2}(r)\|A\|$ (5.36)
for all $A\in\Lambda_{\pi,-l}\cap B(r)^{c}$.
Since $\tilde{\psi}^{\phi}=\omega_{0}$ is a homogeneous product state it now follows from Lemma A.1 that the states $\tilde{\omega}^{\phi}$ are all $g$-close to $\omega_{0}$ for $\phi\in[0,\pi]$ for a $g\in{\mathcal{F}}$ that depends only on $g_{1},g_{2}$, and by Proposition A.2 there is an $\tilde{f}\in{\mathcal{F}}$ depending only on $g$, and $\tilde{f}$-local homogeneous unitaries $W_{\phi}\in{\mathcal{A}}^{\mathrm{al}}$ such that $\tilde{\omega}^{\phi}=\omega_{0}\circ\mathrm{Ad}[W_{\phi}]$. It follows that $\omega^{\phi}=\psi^{\phi}\circ\mathrm{Ad}[V_{\phi}]$ with $V_{\phi}=(\alpha\circ\alpha_{\phi}^{K})^{-1}(W_{\phi})$, which are uniformly $f$-local for some $f\in{\mathcal{F}}$.
@@
We have $\mathrm{Ind}_{2}(\omega_{{\mathscr{P}}})=\operatorname{\mathrm{ind}_{\mathrm{FKM}}}({\mathscr{P}})$.